a. Direct materials prices in 2015 are expected to increase by 8% compared with 2014. b. Under the terms of the labor contract, direct manufacturing labor wage rates are expected to increase c. Power rates and wage rates for supervision, materials handling, and maintenance are not expected to d. Depreciation costs are expected to increase by 6%, and rent, property taxes, and administration costs e. Hamwey expects to manufacture and sell 90,000 units in 2015. by 12% in 2015 compared with 2014. change from 2014 to 2015 are expected to increase by 12%. Print none
media%2F831%2F8316704e-c6dd-468f-9948-c2
media%2Fedc%2Fedc327d5-e9e9-4698-9b54-d2
Answer #1
Manufacturing Cost Classification for the Year 2014 as follows:
Account Total Costs VC as % of TC Variable Costs   Fixed Costs Variable Costs per Unit
1 2 3 = 1 X 2 4 = 1 -3 5 = 3 / 80,000 Units
Direct Materials                  300,000.00 100%        300,000.00                                  -                               3.75
Direct Manufacturing Labor                  220,000.00 100%        220,000.00                                  -                               2.75
Power                    20,000.00 100%          20,000.00                                  -                               0.25
Supervision Labor                    64,000.00 25%          16,000.00                  48,000.00                             0.20
Materials-Handling Labor                    70,000.00 40%          28,000.00                  42,000.00                             0.35
Maintenance Labor                    60,000.00 40%          24,000.00                  36,000.00                             0.30
Depreciation                    88,000.00 0%                          -                    88,000.00                                 -  
Rent, Property Taxes and Administration                  105,000.00 0%                          -                  105,000.00                                 -  
Total                  927,000.00        608,000.00                319,000.00
Total Manufacturing Cost for 2014 = $927,000
Variable Costs for 2015:
Variable Costs per Unit for 2014 Percentage Increase Increase in per Unit Variable Cost per unit for 2015 Total Variable Costs for 2015
5 6 7 = 5 X 6 8 = 5 + 7 9 = 8 X 90,000
Direct Materials                               3.75 8%                     0.30                             4.05                364,500.00
Direct Manufacturing Labor                               2.75 12%                     0.33                             3.08                277,200.00
Power                               0.25                      -                            -                               0.25                  22,500.00
Supervision Labor                               0.20                      -                            -                               0.20                  18,000.00
Materials-Handling Labor                               0.35                      -                            -                               0.35                  31,500.00
Maintenance Labor                               0.30                      -                            -                               0.30                  27,000.00
Depreciation                                   -                        -                            -                                    -                                   -  
Rent, Property Taxes and Administration                                   -                        -                            -                                    -                                   -  
Total                740,700.00
Fixed and Total Costs in 2015
Fixed Costs for 2014 Percentage Increase Increase in FC Fixed Costs for 2015 Variable Cost for 2015 Total Costs
10 11 12 = 10 X 11 13 = 10 + 12 14 15 = 13 + 14
Direct Materials                                   -                        -                            -                                    -                  364,500.00         364,500.00
Direct Manufacturing Labor                                   -                        -                            -                                    -                  277,200.00         277,200.00
Power                                   -                        -                            -                                    -                    22,500.00           22,500.00
Supervision Labor                    48,000.00                      -                            -                    48,000.00                  18,000.00           66,000.00
Materials-Handling Labor                    42,000.00                      -                            -                    42,000.00                  31,500.00           73,500.00
Maintenance Labor                    36,000.00                      -                            -                    36,000.00                  27,000.00           63,000.00
Depreciation                    88,000.00 6%             5,280.00                  93,280.00                                 -             93,280.00
Rent, Property Taxes and Administration                  105,000.00 12%          12,600.00                117,600.00                                 -           117,600.00
Total                  319,000.00          17,880.00                336,880.00                740,700.00     1,077,580.00
Total Manufacturing Cost for 2015 = $1,077,580